OpenStudy (anonymous):

what is the vertex of the equation y=x^2-4x+4

OpenStudy (yttrium):

Try to convert it into it's standard form.

OpenStudy (yttrium):

\[y-k = \pm4a (x-h)^2\]

OpenStudy (yttrium):

In that case. It is \[y-0 = (x-2)^2\]

OpenStudy (yttrium):

Therefore, the vertex is at point. (h,k)

OpenStudy (yttrium):

So what is that @Jrigby20 ??

OpenStudy (anonymous):

(2,0)?

OpenStudy (yttrium):

you got this

OpenStudy (anonymous):

so thats right?

OpenStudy (yttrium):

yea

OpenStudy (anonymous):

k thanks a ton